OpenStudy (anonymous):

please help me for lab6

OpenStudy (anonymous):

1. Run Tran: Final output voltage is 193mV So, for this voltage divider circuit, 0.193 = VS*RON/(RON+RL). Solving, RON=688 ohms 2. Vth = VS*RON/(RON+RL) = 0.193V Rth = RON || RL = 643.2 ohms 3. Substituting in Eqn 10.66, t = 0.5ns 4. Measured tpd rising is 0.6ns 5. Measured tpd falling is 3.45ns 6. Run Tran: Measure cycle time: approx (23.75-7.625)=16.125 Dividing by 9, we get 1.79ns

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

thanks sam